Tech Staff Meeting Notes April 18, 2006

Current Top Priorities are:Station Maintenance, Broadband Upgrade, NC Node

There are four levels of priority.
PrioritiesWish to try to get these done (or make progress on) before the next meeting (yeah right)
Second TierFit this in too if an opportunity arises. May be good to do instead of priorities but check with Mitch. Likely to be a priority in a previous and/or subsequent month
As Time PermitsFit these in when you can but not instead of first or second priorities
OtherStuff we hope to get to and don't want to forget about

This redistributed task list agreed to at the November meeting. Note that everyone is still needed to work on essentially all tasks. There is however, a single person who is primarily responsible for each one. Mitch of course, is ultimately responsible for everything.

Greg's March ToDo List
is only partially incorporated into this rest of this page.

This meeting focused only on the priorities, the rest of the notes are carried forward. Let me know if there is anything to add to things not covered in the meeting.


BB network upgrade.
Huddle tests are in progress and several sensors were blessed by Greg. See Mitch's PDF webpage for results.
MKTA node will be upgraded beginning this week. Need a sensor asap to replace GNAR (rather than swap a dm16 for a dm16 then a dm24).
NMAD node is after that (need telem plan).
FIFO mode in the dm24's will be configured to overwrite oldest (as opposed to write once). If we lose comm, this will cause loss of the oldest data. The thinking is these are broadband stations not strongmo so if given the Solomonic choice, prefer to sacrifice the oldest to save the youngest.
Be sure to contact Mitch, Steve, or Al when a when doing a specific station so mods to the earthworm and to the instrument response files can be modified appropriately. We're going from 6 channels to 3 and the digitizer constant will change.

Bob was at this meeting. James will be hitting 4 station upgrades beginning the end of the month (after classes and thesis revisions). The rest of these notes are carried from before.
The GPS Project needs a maintenace visit to CJTR at Camp Robinson in North Litte Rock, Arkansas. It's been a year since the last visit. The battery, enclosure seals, fans and filters, and timer battery need to be checked and/or replaced. In addition, we need updated contact info. Specifically, we need to know whom to contact regarding getting CJTR on the Internet.

Notes from previous meeting are carried along from January.
Columbia, TN permission is a wash. Still don't have an eastern hard rock station.
Need to notify James when we do the fiber isolation at the CVTN strongmo station so he can verify we didn't muck up the GPS station.
Camp Robinson needs a routine service. Contact info needs to be updated and is likely best done on site.
Need to test aerocom telemetry path at Steele. Station is at private business near airport and the plan is to telem to High School on the other side of town where we can get internet.
Piggot not yet converted to ethernet though permission is secured. IT tech is difficulty to contact.
Need a new plan for MS County Airport at East Prairie, MO.
Reelfoot needs to be converted from Solar to AC.

ETN The map shows stations with yellow triangles being new ones.
John is here with the former LRNC node. Chris will meet John the week of May 8 to rehab as many NC stations as possible. Greg, Jim, Bill, and Al will meet Chris and John the following week (May 15) to install the new node at Mt. Mitchell. Al and Steve are working on the PC to make it as plug and play as possible.

Need to swap SEAR with 2g. Done.
Continued problems with mss100's losing connectivity (SMOC: simple matter of connectivity). The good news is Jim verified that the stations are still working when SMOC is lost.
SWET-3H52, ADSC-5657, and CSTN-5637 need to have the leap second corrected. SHTN-5648 and TRSC-5662 might also but need a site visit to verify.
Need to upgrade trimble firmware to avoid leap second error next time. A workable upgrade procedure doesn't yet exist. Jim, Wurth, and DTA or working on it. Done.
CVTN and ASAR optical isolation. Need to upgrade to year 3 hardware. Hardware ready. James brought back some nice pictures from CVTN to help figure out how to do it (e.g. the conduit is way too small).

Routine Maintenance. Jim knows what to do.

Bix will need either Chris or Bill to instruct and verify proper operation of the pneumatic seismic source. Should only be a day maybe two (or part of a day or two). First instance is sometime April 24-28. And second instance sometime early in July.

Second Tier Priorities

Enclosure for external HVAC pieces on the east of 3904 was started and stalled. Need to finish this up and get rid of the two parking spaces worth of gravel.

Need to follow up on ATV training and helmet(s).

NCCU Gordana Vlahovic is getting a cmg6td to install at North Carolina Central University. John worked out a plan with her. Greg worked up a hardware cost estimate and Mitch faxed it so they can generate a PO.

SWET Upgrade This has gone no where since last time. Bran Potter sent a list of power options. Alena gave some specs for solar operation. Looks like the best way is to do a low voltage feed like TZTN. Should we use VSAT? Still need to somehow provide real-time wiggle display in the Geology building.

Steiner Brain Dump Seminar Beginning this summer, Greg will be giving a weekly non-credit seminar. It will primarily be for the tech staff and will cover all the knowledge we need to pass from Greg to other staff. Others at CERI will be welcome to participate if they wish. Preferred time is Mondays at 9. This will start sometime in mid-May. Send ideas on what to want to learn about to Greg.

3904/211 New EW and Hydra responsibilities require more space for testing and configuring computer hardware. This new space will be in 3904/211. The copier will go where the light table is and the light table will go where ever JM Chiu wants it to. Done.

Mizzou Brent Rosenblad at UofMO, Columbia will be doing surface wave analysis with a large vibroseis source in the NMSZ. Jim or Chris will meet Brent in Ridgely tuesday, April 4 at 9:30 to tour a couple stations and work out a plan.

FDL. Mitch needs to yank Paul's chain to get an account number for new laptops. He also needs to yank the chains of Steve and Paul to get a list of required software.

As Time Permits

I-40 Bridge.
This project is in O&M mode--no more development. Al will need to clean off the recorder disks every other week. We can't afford to spend anymore time improving at the expense of other projects. Hence we will attempt to keep it going as best we can. Doesn't work too bad given all the possible points of failure.


Got the modem and internet works. Need to integrate into CERI central processing.

MPHB. Withers needs to determine current orientation so we can pull it and orient correctly. Helicorders look strange. Need to verify the data.

VBMS. Vault installation went out for bid.

Langston student projects continue. See Greg's list above.